# See Less, See Right: Bi-directional Perceptual Shaping For Multimodal Reasoning

### v1

@@ 1 Introduction
Large vision–language models (VLMs) are increasingly serving as a unified interface for both visual and language-based reasoning [bai2025qwen2, chen2024internvl]. Among real-world applications, visual question answering (VQA) is a widely deployed, high-impact task: a system must parse a natural-language query, localize the pertinent visual evidence, and produce an answer whose reasoning remains tied to that evidence. Despite rapid progress, the perceptual capability of VLMs, including the identification, localization, and accurate reading of fine-grained visual cues is still a bottleneck [liu2025perception, sim2025can, bigverdi2025perception]. If perception slips, downstream reasoning can rely on incomplete or misleading cues, yielding plausible but evidence-mismatched answers and degrading VQA performance.
@@
To mitigate this perceptual bottleneck, a complementary line of work augments VLMs with external visual tools (e.g., cropping, masking, segmentation) that produce evidence-focused intermediate visual cues at inference time [shao2024visual, wang2025multimodal, fu2025refocus, zhang2025pixelcraft, zheng2024instruction, hu2024visual, man2025argus]. Recent efforts further collect step-by-step “visual chain-of-thought” traces, where the model is guided by intermediate boxes, tool-use trajectories, or auxiliary images and then trained to reproduce these visual reasoning steps [shao2024visual, Chen2025MINTCoTEI, fu2025refocus, zheng2024instruction, yang2025machinementalimageryempower]. In practice, such mechanisms improve grounding and answer accuracy across tasks such as chart understanding, image-based math problems, and natural-image VQA with sparse cues.
@@
In this paper, we propose Bi-directional Perceptual Shaping (BiPS), a two-stage training-time approach integrated into the Group Relative Policy Optimization (GRPO) framework [shao2024grpo], which shapes a VLM’s perception by pulling predictions toward an evidence-preserving view and pushing them away from an evidence-ablated view. BiPS first builds a question-conditioned evidence-preserving view that keeps regions needed to answer the query while masking distractors, and applies a consistency constraint based on the Kullback–Leibler (KL) divergence to align the model’s predictions on this view with those on the original image, encouraging coarse but complete coverage of supporting pixels. It then constructs a complementary evidence-ablated view that finely removes critical pixels so the answer changes or becomes unanswerable, and adds a KL-based separation term that pushes predictions on the original and ablated views apart, discouraging text-only shortcuts (i.e., answering from text alone) and enforcing fine-grained visual reliance. This bidirectional shaping yields visually grounded decisions while requiring no extra annotations or customized parsers at inference.
Realizing BiPS requires precise question-conditioned supervision in the form of paired views, including an evidence-preserving view and an evidence-ablated view for each image and question pair. Naive pixel-level masking or cropping is common, but it remains coarse and shape-constrained. Complex, multi-panel charts naturally carry the fine-grained evidence (e.g., thin polylines, layered marks, and small symbols), making charts a rich source of training signals. We therefore build a programmatic data pipeline for chart data that generates the required evidence-preserving and evidence-ablated views. In this work, we instantiate the pipeline using ECD [yang2025ecd10k], a synthetic corpus of complex multi-panel figures paired with executable rendering code. Because each figure is generated by code, every object (marks, layers, axes, legends) has explicit provenance, which enables exact edits to synthesize the two complementary views. This pipeline yields 13K synthetic training examples. Fine-tuned solely on this set, BiPS already generalizes well: across eight benchmarks spanning real-world figure datasets (e.g., CharXiv [wang2024charxiv], ChartQAPro [masry2025chartqapro]) and out-of-domain general VQA (e.g., MathVista [lu2023mathvista], MMStar [chen2024mmstar]), average accuracy improves by +7.3% over the base model (Qwen2.5-VL-7B [bai2025qwen2]). Adding 39k math-specialized examples with standard GRPO further increases the average gain to 8.2%.
@@ 2.1 Multimodal Reasoning
Recent advances in reinforcement learning (RL) for language models have sparked growing efforts to extend RL to multimodal reasoning [Luo2025Ursa, zhang2025generative]. Recent studies apply RL to strengthen visual understanding or reasoning across domains, including single-image [meng2025mmeureka, yang2025r1onevision, huang2025visionr1, Chen2025revisualr1, wang2025vlrethinker] reasoning for spatial perception, multi-image reasoning [zhang2025perl] for cross-scene consistency, video reasoning  [Feng2025VideoR1RV] for temporal causality, and chart reasoning  [chen2025chartr1] for quantitative alignment.
Complementary to advances in language reasoning, a growing body of work improves VLMs by injecting intermediate visual cues at inference time. Typical strategies highlight question-relevant content via bounding boxes [shao2024visual], question-conditioned crops [man2025argus], or masks that suppress distractors [zheng2024instruction], biasing predictions toward the appropriate visual evidence. Building on this idea, subsequent methods design task-aware tools or specialized modules for tighter evidence localization, such as chart-specific annotation tools [fu2025refocus, zhang2025pixelcraft] and visual sketchpad [hu2024visual]. Pushing this paradigm further, recent studies supervise models with labeled evidence and train them to reproduce coordinates, masks, or tool selections so that downstream answers can be conditioned on the predicted evidence [fu2025refocus, shao2024visual, zheng2024instruction, su2025openthinkimg, wu2025vtool, sarch2025grounded, Zheng2025DeepEyesI, Su2025PixelReasoner, Chen2025MINTCoTEI]. However, these designs inherit several limitations: rectangular or coarse masks struggle with fine-grained or irregular structures, task-specific engineering reduces generalization, and multi-step pipelines introduce nontrivial inference overhead.
A recent work removes the need for explicit intermediate images by encouraging models to reason through latent internal processes [yang2025machinementalimageryempower], but such approach remain confined to specific tasks. ChiP [fu2025chip] and PAPO [wang2025papo] inject Gaussian noise or random masks as negative perturbations that penalize reliance on incorrect images to avoid visual hallucinations and text-only shortcuts, yet these methods overlook informative detailed signals inside the figure.
@@ 2.2 Chart Understanding And Reasoning
Unlike general multimodal reasoning on natural scenes [Fu2023MMEAC] or geometric diagrams [Lu2021InterGPSIG, lu2023mathvista, wang2024mathvision], chart understanding targets structured quantitative graphics where numerical relations are encoded by axes and marks, demanding precise value perception and visual–numerical correspondence. Early chart-domain studies mainly focused on low-level perception tasks such as chart element detection and text extraction [Methani2019PlotQARO, Masry2022ChartQAAB], while recent benchmarks shift toward reasoning-centric tasks [wang2024charxiv, tang2025chartmuseum, masry2025chartqapro] that require interpreting implicit patterns, making comparisons, and executing multi-step computations over visualized data. To enhance such reasoning, subsequent studies convert charts into structured symbolic programs or code for executable reasoning [Zhang2024TinyChartEC, Jia2025ChartReasonerCM, Zhao2025ChartCoderAM, yang2024chartmimic], while others explore multimodal feedback, reflective learning, and reinforcement-based optimization [Huang2025ChartSketcherRW, masry2025bigcharts, chen2025chartr1] to improve reliability. Nonetheless, chart reasoning still faces challenges such as limited chart-to-code accuracy [Masry2022ChartQAAB, Xu2023ChartBenchAB], weak generalization to diverse layouts, and difficulty handling high-frequency curve fluctuations or subtle oscillatory patterns, which often cause quantitative misalignment. Addressing these challenges demands finer-grained perception and robust visual reasoning.
@@ 3 Method
To address the core perceptual failures of VLMs, namely their tendency to be distracted by irrelevant visual information and their inability to focus on fine-grained evidence, we propose BiPS. As illustrated in Figure 2, BiPS is implemented as a two-stage training curriculum consisting of a Consistency Stage followed by a Separation Stage. The Consistency Stage guides the model to maintain consistent predictions when only evidence-relevant regions are retained, whereas the Separation Stage drives the model to diverge when the critical regions are removed. Both stages are implemented as bidirectional KL-based objectives, optimized within the GRPO [shao2024grpo] framework.
@@ 3.1 Programmatic Data Construction Pipeline
We overcome this limitation by constructing the paired views programmatically in the code domain. Our data generation pipeline builds upon ECD [yang2025ecd10k], which provides complex, multi-panel charts along with their executable rendering code. Charts serve as an ideal substrate for this purpose: their elements (marks, axes, legends) are semantically structured, fine-grained, and explicitly linked to the underlying code. This enables editing at the code level rather than pixels, yielding semantically precise and fine-grained control over visual evidence. As shown in Figure 3, our pipeline consists of the following three main stages.
@@
The original dataset’s reasoning questions are open-ended and difficult to verify by predefined rules. Following a typical RLVR setting [shao2024grpo, meng2025mmeureka], we refine this data by employing an auxiliary LLM arbitrator, GPT5-mini, to convert the original questions into a multiple-choice format. This arbitrator is provided with the chart’s source code and metadata to ensure that the reformulated question remains answerable and that the ground-truth option is correct. This step provides verifiable supervision compatible with downstream RL training and ensures the overall quality of our base question set.
@@
To focus the training on non-trivial examples, we filter out questions that are too “easy” for the base model (Qwen2.5-VL-7B-Instruct [bai2025qwen2]). We perform 8 rollouts for each validated question. Any question that the base model answers correctly in all rollouts is considered “easy” and is discarded from our training set, allowing our method to concentrate on more challenging reasoning tasks.
@@ 3.2.2 Robustness via Separation
On its own, $\mathcal{L}_{\text{cons}}$ is insufficient as it is susceptible to “shortcut learning”. A model can perfectly minimize $\mathcal{L}_{\text{cons}}$ by ignoring both visual inputs and relying solely on the textual question $q$ to produce the same answer.
@@ 3.3 A Coarse-to-Fine Training Curriculum
Before detailing the stages, we define the base GRPO [shao2024grpo] objective, which extends Proximal Policy Optimization (PPO) by normalizing rewards across rollouts within the same group to stabilize training. The training objective is:
@@ 4 Experiments
Table 1: Evaluation on chart understanding and general perception & reasoning benchmarks. Models with [Uncaptioned image] are chart-specialized; models with [Uncaptioned image] are math-specialized; denotes our models. Blue numbers denote chart-related data, and red numbers denote math-related or perception-related data. Avg. is the arithmetic mean over available metrics. Models trained on MathVision (marked with ) omit their scores on MathVision. Best scores are in bold, and second-best scores are underlined.
@@ 4.1 Experiment Setting
We evaluate our models on a comprehensive suite of benchmarks spanning two categories. To assess chart understanding and reasoning capabilities, we report scores on CharXiv [wang2024charxiv], ChartQAPro [masry2025chartqapro], ChartMuseum [tang2025chartmuseum], Evochart [huang2025evochart] and ECD-Bench [yang2025ecd10k]. For general perception and reasoning, We use MathVista [lu2023mathvista], MathVision [wang2024mathvision], MathVerse-VO [zhang2024mathverse], and MMStar [chen2024mmstar].
@@
We evaluate our method against diverse baselines, including closed-source models like GPT-4o [hurst2024gpt] and Claude-3.7-Sonnet as high-level SOTA references. Our open-source comparisons include general models such as InternVL3-8B [zhu2025internvl3] and Qwen2.5-VL-7B [bai2025qwen2]. We also compare against specialized reasoning models, including chart-focused systems (ChartLlama [han2023chartllama], ChartGemma [masry2025chartgemma], Chart-R1 [chen2025chartr1], BigCharts-R1 [masry2025bigcharts]) and broader multimodal reasoners (R1-OneVision [yang2025r1onevision], Vision-R1 [huang2025visionr1], DeepEyes [Zheng2025DeepEyesI]) that handle both chart and math reasoning.
@@ 4.3 Ablation Study
We compare our programmatic code-editing strategy against a random masking baseline. Following prior work [wang2025papo, wang2024mdpo], this baseline randomly masks out 60% of the image patches to create the alternative view. As shown in Table 4, our programmatic code-editing strategy significantly outperforms the random masking baseline across all benchmarks. These improvements highlight the importance of generating semantically faithful counterparts for the bi-directional KL constraints. Programmatic editing explicitly isolates the key-evidence and question-unrelated code components, ensuring that $\mathcal{L}_{\text{sep}}$ is computed on a pure distraction view. In contrast, random masking drops patches blindly and fails to produce a clean distraction view: if it mostly hides irrelevant regions, the example becomes easier and the prediction gets closer to $I_{\text{pres}}$, making KL maximization counterproductive; when it also hides task-critical evidence, the masked input becomes a mixture of missing-evidence and residual distractions rather than a pure distractor, so the KL is computed against an ill-defined target and yields noisy gradients. By ensuring clean, semantically controlled counterparts, our programmatic generation provides precise supervision for both $\mathcal{L}_{\text{cons}}$ and $\mathcal{L}_{\text{sep}}$, leading to more stable optimization and stronger perceptual grounding.
@@
Figure 5 presents two representative cases from ChartXiv [wang2024charxiv]. In both cases, Qwen2.5-VL-7B produces plausible but incorrect answers by relying on textual or statistical cues rather than interpreting the visual structure. In the intersection-count example, it hallucinates extra curve crossings without actually tracing the plotted functions, while in the multi-plot accuracy example it overfits to frequently occurring numerical patterns (e.g., 94.6%) instead of reasoning over per-panel maxima. In contrast, BiPS yields correct and visually grounded responses. This improvement suggests the bidirectional perceptual shaping encourages the model to attend to the structural alignment and cross-plot relationships that are critical for reasoning, rather than inferring from superficial patterns or numeric priors. Additional cases can be found in the Appendix.

### v2

@@ 1 Introduction
Large vision–language models (VLMs) are increasingly serving as a unified interface for both visual and language-based reasoning [1, 7]. Among real-world applications, visual question answering (VQA) is a widely deployed, high-impact task: a system must parse a natural-language query, localize the pertinent visual evidence, and produce an answer whose reasoning remains tied to that evidence. Despite rapid progress, the perceptual capability of VLMs, including the identification, localization, and accurate reading of fine-grained visual cues is still a bottleneck [20, 35, 2]. If perception slips, downstream reasoning can rely on incomplete or misleading cues, yielding plausible but evidence-mismatched answers and degrading VQA performance.
@@
To mitigate this perceptual bottleneck, a complementary line of work augments VLMs with external visual tools (e.g., cropping, masking, segmentation) that produce evidence-focused intermediate visual cues at inference time [33, 42, 11, 53, 57, 13, 24]. Recent efforts further collect step-by-step “visual chain-of-thought” traces, where the model is guided by intermediate boxes, tool-use trajectories, or auxiliary images and then trained to reproduce these visual reasoning steps [33, 6, 11, 57, 50]. In practice, such mechanisms improve grounding and answer accuracy across tasks such as chart understanding, image-based math problems, and natural-image VQA with sparse cues.
@@
In this paper, we propose Bi-directional Perceptual Shaping (BiPS), a two-stage training-time approach integrated into the Group Relative Policy Optimization (GRPO) framework [34], which shapes a VLM’s perception by pulling predictions toward an evidence-preserving view and pushing them away from an evidence-ablated view. BiPS first builds a question-conditioned evidence-preserving view that keeps regions needed to answer the query while masking distractors, and applies a consistency constraint based on the Kullback–Leibler (KL) divergence to align the model’s predictions on this view with those on the original image, encouraging coarse but complete coverage of supporting pixels. It then constructs a complementary evidence-ablated view that finely removes critical pixels so the answer changes or becomes unanswerable, and adds a KL-based separation term that pushes predictions on the original and ablated views apart, discouraging text-only shortcuts (i.e., answering from text alone) and enforcing fine-grained visual reliance. This bidirectional shaping yields visually grounded decisions while requiring no extra annotations or customized parsers at inference.
Realizing BiPS requires precise question-conditioned supervision in the form of paired views, including an evidence-preserving view and an evidence-ablated view for each image and question pair. Naive pixel-level masking or cropping is common, but it remains coarse and shape-constrained. Complex, multi-panel charts naturally carry the fine-grained evidence (e.g., thin polylines, layered marks, and small symbols), making charts a rich source of training signals. We therefore build a programmatic data pipeline for chart data that generates the required evidence-preserving and evidence-ablated views. In this work, we instantiate the pipeline using ECD [49], a synthetic corpus of complex multi-panel figures paired with executable rendering code. Because each figure is generated by code, every object (marks, layers, axes, legends) has explicit provenance, which enables exact edits to synthesize the two complementary views. This pipeline yields 13K synthetic training examples. Fine-tuned solely on this set, BiPS already generalizes well: across eight benchmarks spanning real-world figure datasets (e.g., CharXiv [44], ChartQAPro [25]) and out-of-domain general VQA (e.g., MathVista [21], MMStar [4]), average accuracy improves by +7.3% over the base model (Qwen2.5-VL-7B [1]). Adding 39k math-specialized examples with standard GRPO further increases the average gain to 8.2%.
@@ 2.1 Multimodal Reasoning
Recent advances in reinforcement learning (RL) for language models have sparked growing efforts to extend RL to multimodal reasoning [23, 54]. Recent studies apply RL to strengthen visual understanding or reasoning across domains, including single-image [29, 48, 16, 5, 40] reasoning for spatial perception, multi-image reasoning [55] for cross-scene consistency, video reasoning  [8] for temporal causality, and chart reasoning  [3] for quantitative alignment.
Complementary to advances in language reasoning, a growing body of work improves VLMs by injecting intermediate visual cues at inference time. Typical strategies highlight question-relevant content via bounding boxes [33], question-conditioned crops [24], or masks that suppress distractors [57], biasing predictions toward the appropriate visual evidence. Building on this idea, subsequent methods design task-aware tools or specialized modules for tighter evidence localization, such as chart-specific annotation tools [11, 53] and visual sketchpad [13]. Pushing this paradigm further, recent studies supervise models with labeled evidence and train them to reproduce coordinates, masks, or tool selections so that downstream answers can be conditioned on the predicted evidence [11, 33, 57, 37, 45, 32, 58, 36, 6]. However, these designs inherit several limitations: rectangular or coarse masks struggle with fine-grained or irregular structures, task-specific engineering reduces generalization, and multi-step pipelines introduce nontrivial inference overhead.
A recent work removes the need for explicit intermediate images by encouraging models to reason through latent internal processes [50], but such approach remain confined to specific tasks. ChiP [10] and PAPO [43] inject Gaussian noise or random masks as negative perturbations that penalize reliance on incorrect images to avoid visual hallucinations and text-only shortcuts, yet these methods overlook informative detailed signals inside the figure.
@@ 2.2 Chart Understanding And Reasoning
Unlike general multimodal reasoning on natural scenes [9] or geometric diagrams [22, 21, 41], chart understanding targets structured quantitative graphics where numerical relations are encoded by axes and marks, demanding precise value perception and visual–numerical correspondence. Early chart-domain studies mainly focused on low-level perception tasks such as chart element detection and text extraction [30, 26], while recent benchmarks shift toward reasoning-centric tasks [44, 38, 25] that require interpreting implicit patterns, making comparisons, and executing multi-step computations over visualized data. To enhance such reasoning, subsequent studies convert charts into structured symbolic programs or code for executable reasoning [51, 18, 56, 47], while others explore multimodal feedback, reflective learning, and reinforcement-based optimization [15, 27, 3] to improve reliability. Nonetheless, chart reasoning still faces challenges such as limited chart-to-code accuracy [26, 46], weak generalization to diverse layouts, and difficulty handling high-frequency curve fluctuations or subtle oscillatory patterns, which often cause quantitative misalignment. Addressing these challenges demands finer-grained perception and robust visual reasoning.
@@ 3 Method
To address the core perceptual failures of VLMs, namely their tendency to be distracted by irrelevant visual information and their inability to focus on fine-grained evidence, we propose BiPS. As illustrated in Figure 2, BiPS is implemented as a two-stage training curriculum consisting of a Consistency Stage followed by a Separation Stage. The Consistency Stage guides the model to maintain consistent predictions when only evidence-relevant regions are retained, whereas the Separation Stage drives the model to diverge when the critical regions are removed. Both stages are implemented as bidirectional KL-based objectives, optimized within the GRPO [34] framework.
@@ 3.1 Programmatic Data Construction Pipeline
We overcome this limitation by constructing the paired views programmatically in the code domain. Our data generation pipeline builds upon ECD [49], which provides complex, multi-panel charts along with their executable rendering code. Charts serve as an ideal substrate for this purpose: their elements (marks, axes, legends) are semantically structured, fine-grained, and explicitly linked to the underlying code. This enables editing at the code level rather than pixels, yielding semantically precise and fine-grained control over visual evidence. As shown in Figure 3, our pipeline consists of the following three main stages.
@@
The original dataset’s reasoning questions are open-ended and difficult to verify by predefined rules. Following a typical RLVR setting [34, 29], we refine this data by employing an auxiliary LLM arbitrator, GPT5-mini, to convert the original questions into a multiple-choice format. This arbitrator is provided with the chart’s source code and metadata to ensure that the reformulated question remains answerable and that the ground-truth option is correct. This step provides verifiable supervision compatible with downstream RL training and ensures the overall quality of our base question set.
@@
To focus the training on non-trivial examples, we filter out questions that are too “easy” for the base model (Qwen2.5-VL-7B-Instruct [1]). We perform 8 rollouts for each validated question. Any question that the base model answers correctly in all rollouts is considered “easy” and is discarded from our training set, allowing our method to concentrate on more challenging reasoning tasks.
@@ 3.2.2 Robustness via Separation
On its own, $\mathcal{L}_{\text{cons}}$ is insufficient as it is susceptible to “shortcut learning”. A model may satisfy the consistency $\mathcal{L}_{\text{cons}}$ by exploiting surrounding text (OCR) or language priors to produce identical answers for the original and evidence-preserving images, without attending to fine-grained details (e.g., curves).
@@ 3.3 A Coarse-to-Fine Training Curriculum
Before detailing the stages, we define the base GRPO [34] objective, which extends Proximal Policy Optimization (PPO) by normalizing rewards across rollouts within the same group to stabilize training. The training objective is:
@@ 4 Experiments
Table 1: Evaluation on chart understanding and general perception & reasoning benchmarks. Models with [Uncaptioned image] are chart-specialized; models with [Uncaptioned image] are math-specialized; denotes our models. Blue numbers denote chart-related data, and red numbers denote math-related or perception-related data. Avg. is the arithmetic mean over available metrics. Models trained on MathVision (marked with ) omit their scores on MathVision. Best scores are in bold, and second-best scores are underlined. Baselines are reproduced using VLMEvalKit to ensure strict fairness.
@@ 4.1 Experiment Setting
We evaluate our models on a comprehensive suite of benchmarks spanning two categories. To assess chart understanding and reasoning capabilities, we report scores on CharXiv [44], ChartQAPro [25], ChartMuseum [38], Evochart [14] and ECD-Bench [49]. For general perception and reasoning, We use MathVista [21], MathVision [41], MathVerse-VO [52], and MMStar [4].
@@
We evaluate our method against diverse baselines, including closed-source models like GPT-4o [17] and Claude-3.7-Sonnet as high-level SOTA references. Our open-source comparisons include general models such as InternVL3-8B [59] and Qwen2.5-VL-7B [1]. We also compare against specialized reasoning models, including chart-focused systems (ChartLlama [12], ChartGemma [28], Chart-R1 [3], BigCharts-R1 [27]) and broader multimodal reasoners (R1-OneVision [48], Vision-R1 [16], DeepEyes [58]) that handle both chart and math reasoning.
@@ 4.2 Main Results
To verify that the improvements stem from our specific perceptual shaping pipeline rather than merely applying RL to the data, we compare BiPS-General-7B against a baseline trained with standard GRPO. This baseline is fine-tuned on the exact same combined dataset (programmatic samples + ViRL39k) but treats all data uniformly without the proposed two-stage perceptual shaping curriculum. While standard GRPO yields substantial gains over the base model, BiPS-General consistently outperforms it across all benchmarks. Notably, on complex chart reasoning tasks like CharXiv, our method surpasses the standard GRPO baseline by a significant margin (+5.2). This performance gap highlights a critical insight: simply optimizing reasoning via RL is insufficient if the underlying visual grounding is flawed. By explicitly shaping the model’s perception through our programmatic curriculum, BiPS ensures that the RL process operates on high-fidelity visual signals, thereby amplifying the effectiveness of the optimization.
@@ 4.3 Ablation Study
We compare our programmatic code-editing strategy against a random masking baseline. Following prior work [43, 39], this baseline randomly masks out 60% of the image patches to create the alternative view. As shown in Table 4, our programmatic code-editing strategy significantly outperforms the random masking baseline across all benchmarks. These improvements highlight the importance of generating semantically faithful counterparts for the bi-directional KL constraints. Programmatic editing explicitly isolates the key-evidence and question-unrelated code components, ensuring that $\mathcal{L}_{\text{sep}}$ is computed on a pure distraction view. In contrast, random masking drops patches blindly and fails to produce a clean distraction view: if it mostly hides irrelevant regions, the example becomes easier and the prediction gets closer to $I_{\text{pres}}$, making KL maximization counterproductive; when it also hides task-critical evidence, the masked input becomes a mixture of missing-evidence and residual distractions rather than a pure distractor, so the KL is computed against an ill-defined target and yields noisy gradients. By ensuring clean, semantically controlled counterparts, our programmatic generation provides precise supervision for both $\mathcal{L}_{\text{cons}}$ and $\mathcal{L}_{\text{sep}}$, leading to more stable optimization and stronger perceptual grounding.
@@
Figure 5 presents two representative cases from ChartXiv [44]. In both cases, Qwen2.5-VL-7B produces plausible but incorrect answers by relying on textual or statistical cues rather than interpreting the visual structure. In the intersection-count example, it hallucinates extra curve crossings without actually tracing the plotted functions, while in the multi-plot accuracy example it overfits to frequently occurring numerical patterns (e.g., 94.6%) instead of reasoning over per-panel maxima. In contrast, BiPS yields correct and visually grounded responses. This improvement suggests the bidirectional perceptual shaping encourages the model to attend to the structural alignment and cross-plot relationships that are critical for reasoning, rather than inferring from superficial patterns or numeric priors. Additional cases can be found in the Appendix.